Suppose that Charlie makes a free throw 0.82 on any one try. Assuming that this probability doesn’t change, find the chance that Charlie makes 4 out of the next seven free throws. Find the mean and standard deviation


1
Expert's answer
2021-12-14T13:04:52-0500

Let "X=" the number of successful throws: "X\sim Bin(n, p)."

Given "n=7, p=0.82, q=1-p=0.18."

i)


"P(X=4)=\dbinom{7}{4}(0.82)^4 (0.18)^{7-4}"

"=0.0922870936512"

ii)


"E(X)=np=7(0.82)=5.74"

"Var(x)=\sigma^2=npq=7(0.82)(0.18)=1.0332"

"\sigma=\sqrt{\sigma^2}=\sqrt{1.0332}\approx1.016464"